# HBAR entropy in causal diamond geometry: A near-horizon perspective

### v1

@@ IV.3 Thermal behavior: radiation field density matrix and master equation
which matches the form expected for a thermal state satisfying detailed balance, governed by the Boltzmann factor:
@@
where the index $j$ labels the field modes $\mathbf{s}_{j}$. The diagonal density matrix, written in the occupation number basis $\quantity{n}\equiv\quantity{n_{1},n_{2},\cdots,n_{j},\cdots}$, is given by $\rho_{\text{diag}}(\quantity{n})=\rho_{n_{1},n_{2},\cdots;\,n_{1},n_{2},\cdots}$, where $n_{j}\equiv n_{\mathbf{s}_{j}}$ is the occupation number of scalar quanta in the mode $\mathbf{s}_{j}$. The notation $\quantity{n}_{n_{j}+q}\equiv\quantity{n_{1},n_{2},\cdots,n_{j}+q,\cdots}$ denotes a shift of $q\in\mathbb{Z}$ in the occupation number of the $j$-th mode, with all other occupation numbers unchanged.
@@
Because the atoms are injected randomly and the modes evolve independently, the full multimode steady-state density matrix factorizes as a product over single-mode contributions:
@@
from which we obtain the explicit thermal form:
@@
Therefore, the steady-state density matrix of the radiation field corresponds to a thermal distribution at the causal diamond temperature $T_{D}=1/(\pi\alpha)$, in agreement with the results of Eqs. (50)–(52). It is important to note that this density matrix reproduces the expected Planckian form for the steady-state average occupation number:
@@ VI Discussion
In this article, by analyzing the near-horizon region of the causal diamond (CD) spacetime, we identified an emergent conformal symmetry governed by conformal quantum mechanics (CQM), as discussed in Sec. III, which allowed for an analytic treatment of the field modes. Besides elucidating the associated entropy dynamics, this symmetry is essential for computing the emission and absorption spectra of scalar quanta (see Eqs. (48)–(49)) produced and absorbed by the freely falling atomic cloud as it approaches the causal horizon of the CD spacetime. Employing tools from quantum optics, we computed the von Neumann entropy of the radiation field—composed of scalar quanta—and its corresponding flux, given in Eq. (61).
@@ Appendix A Emission rate for an atomic cloud approaching $\mathcal{H}^{+}_{\infty}$
Having established the CQM modes of the field, we can now compute the corresponding emission rate (and, consequently, the absorption rate). To this end, we need to characterize $\eta$ and $\tau$ in terms of $\rho$ in the region $\rho\gg 1$. In our discussion of the geometry in Sec. II.4.2, we found that by performing the substitution $4\tilde{\rho}=\alpha^{2}/\rho$, the near-horizon analysis in the region of interest is simplified in terms of $\tilde{\rho}\ll 1$ (see Eq. (32)). In this parametrization, $\eta$, $\tau$, and the CQM field modes take the form

### v2

@@ IV.3 Thermal behavior: Radiation field density matrix and master equation
which matches the form expected for a thermal state satisfying detailed balance, governed by the Boltzmann factor
@@
where the index $j$ labels the field modes $\mathbf{s}_{j}$. The diagonal density matrix, written in the occupation number basis $\quantity{n}\equiv\quantity{n_{1},n_{2},\cdots,n_{j},\cdots}$, is given by $\rho_{\text{diag}}(\quantity{n})=\rho_{n_{1},n_{2},\cdots;\,n_{1},n_{2},\cdots}$, where $n_{j}\equiv n_{\mathbf{s}_{j}}$ is the occupation number of scalar quanta in the mode $\mathbf{s}_{j}$. The notation $\quantity{n}_{n_{j}+q}\equiv\quantity{n_{1},n_{2},\cdots,n_{j}+q,\cdots}$ denotes a shift of $q\in\mathbb{Z}$ in the occupation number of the $j$th mode, with all other occupation numbers unchanged.
@@
Because the atoms are injected randomly and the modes evolve independently, the full multimode steady-state density matrix factorizes as a product over single-mode contributions
@@
from which we obtain the explicit thermal form
@@
Therefore, the steady-state density matrix of the radiation field corresponds to a thermal distribution at the causal diamond temperature $T_{D}=1/(\pi\alpha)$, in agreement with the results of Eqs. (50)–(52). It is important to note that this density matrix reproduces the expected Planckian form for the steady-state average occupation number
@@ VI Discussion
In this article, by analyzing the near-horizon region of the CD spacetime, we identified an emergent conformal symmetry governed by CQM, as discussed in Sec. III, which allowed for an analytic treatment of the field modes. Besides elucidating the associated entropy dynamics, this symmetry is essential for computing the emission and absorption spectra of scalar quanta [see Eqs. (48)–(49)] produced and absorbed by the freely falling atomic cloud as it approaches the causal horizon of the CD spacetime. Employing tools from quantum optics, we computed the von Neumann entropy of the radiation field—composed of scalar quanta—and its corresponding flux, given in Eq. (61).
@@ Appendix A EMISSION RATE FOR AN ATOMIC CLOUD APPROACHING $\mathcal{H}^{+}_{\infty}$
Having established the CQM modes of the field, we can now compute the corresponding emission rate (and, consequently, the absorption rate). To this end, we need to characterize $\eta$ and $\tau$ in terms of $\rho$ in the region $\rho\gg 1$. In our discussion of the geometry in Sec. II.4.2, we found that by performing the substitution $4\tilde{\rho}=\alpha^{2}/\rho$, the near-horizon analysis in the region of interest is simplified in terms of $\tilde{\rho}\ll 1$ [see Eq. (32)]. In this parametrization, $\eta$, $\tau$, and the CQM field modes take the form